OpenStudy (anonymous):

Solve. (x – 6)^2 = 2

OpenStudy (anonymous):

hint (a-b)^2=a^2-2ab+b^2

OpenStudy (lukebluefive):

Assuming you're solving for x, first try taking the square root of each side.

OpenStudy (anonymous):

(x – 6)^2 = 2 x^2-12x+36=2 add -2 to both sides x^2-12x+34=0

OpenStudy (anonymous):

@Auxuris u made a mistake the third term is 6^2 not 2^2

OpenStudy (anonymous):

in case it wasn't clear, okay basically you have the identity rule of (a^2-2ab+b^2)=(a+b)2 its the second factorisation identity, you've gotta memorise it^^